# Mixed Membership Estimation for Weighted Networks

### v2

@@ 2 The Degree-Corrected Mixed Membership Distribution-Free model
Under DCMMDF, node $i$’s expectation degree is $\sum_{j=1}^{n}\Omega(i,j)=\theta(i)\Pi(i,:)\sum_{j=1}^{n}P\Pi^{\prime}(j,:)$. We see that if there is a node $\bar{i}$ such that $\Pi(\bar{i},:)=\Pi(i,:)$ (i.e., node $\bar{i}$ has the same community membership as node $i$) for $\bar{i}\neq i$, the expectation degree of node $\bar{i}$ may still differ from that of node $i$ if $\theta(\bar{i})\neq\theta(i)$, and this is the reason we call $\theta(i)$ heterogeneity parameter. For visualization, Figure 2 sketches some overlapping (and non-overlapping) networks that can be modeled by our DCMMDF.
@@
By Proposition A.1 of MixedSCORE , DCMMDF is identifiable when Equations (2)-(4) hold, where Proposition A.1 of MixedSCORE has no constraint on $A$’s distribution and it allows $P$ to have negative elements as long as $P$ has unit-diagonals. However, if $P$ has non-unit diagonals, DCMMDF is unidentifiable unless all entries of $\theta$ are the same.
@@
In this remark, we compare our DCMMDF with some previous models.
  • 1.
    When $\mathcal{F}$ is Bernoulli distribution, DCMMDF reduces to DCMM MixedSCORE , i.e., DCMMDF extends DCMM from un-weighted networks to weighted networks and DCMM is a sub-model of DCMMDF.
  • 2.
    When all nodes are pure, DCMMDF reduces to DCDFM DCDFM , i.e., DCMMDF extends DCDFM from non-overlapping weighted networks to overlapping weighted networks.
  • 3.
    When all nodes are pure and all entries of $\theta$ are the same, DCMMDF reduces to DFM qing2021DFM .
  • 4.
    When $\mathcal{F}$ is Bernoulli distribution and all entries of $\theta$ are the same, DCMMDF reduces to MMSB MMSB .
  • 5.
    When $\mathcal{F}$ is Bernoulli distribution and all nodes are pure, DCMMDF reduces to DCSBM DCSBM .
  • 6.
    When $\mathcal{F}$ is Bernoulli distribution, all nodes are pure, and all entries of $\theta$ are the same, DCMMDF reduces to SBM SBM .
  • 7.
    When $\mathcal{F}$ is Bernoulli distribution, $K$ is 1, and all entries of $\theta$ are the same, DCMMDF reduces to the Erdös-Rényi random graph erdos1960evolution .
After designing our model DCMMDF, a random adjacency matrix $A$ with community membership $\Pi$ can be generated from DCMMDF by the following steps.
@@
After generating the random adjacency matrix $A$ by Steps (a)-(c), in this article, we aim at answering the following questions:
@@
    Can we design an algorithm to estimate $\Pi$ with known $A$ and $K$ when $A$ is generated from arbitrary distribution $\mathcal{F}$ satisfying Equation (6) under our DCMMDF?
@@
    Can we design a method to determine the number of communities $K$ for weighted networks generated from DCMMDF?
In the next few sections, we will answer the above questions by applying an efficient spectral algorithm to fit DCMMDF, building a theoretical guarantee of consistency for the algorithm under DCMMDF, and proposing overlapping weighted modularity to infer $K$.
@@ 3 Algorithm
In this section, we aim at answering Q (1) by designing an efficient spectral algorithm to fit DCMMDF. Because $\Omega$ has a low dimensional structure with rank $K$ when $\mathrm{rank}(\Pi)=K,\mathrm{rank}(P)=K$, and all diagonal entries of $\Theta$ are positive, to answer Q1, one natural way to design an algorithm to fit DCMMDF is from the oracle case with known $\Omega$. A good algorithm fitting DCMMDF should satisfy the following nice property: it returns $\Pi$ exactly when using $\Omega$ to replace $A$. Otherwise, if the algorithm can not return $\Pi$ when the input matrix is $\Omega$, it is hard to say that it fits the model well.
@@
To design an algorithm from the oracle case, because $\Omega$ has a low dimensional structure with $K$ nonzero eigenvalues, we let $\Omega=U\Lambda U^{\prime}$ be the eigendecomposition of $\Omega$ such that $U\in\mathbb{R}^{n\times K},\Lambda\in\mathbb{R}^{K\times K}$, and $U^{\prime}U=I_{K}$, where $\Lambda$’s $k$-th diagonal entry is the leading $k$ eigenvalue of $\Omega$ for $k\in[K]$. Let $U_{*}$ be an $n\times K$ matrix such that its $i$-th row is $U_{*}(i,:)=\frac{U(i,:)}{\|U(i,:)\|_{F}}$ for $i\in[n]$. Let $N_{U}$ be an $n\times n$ diagonal matrix such that its $i$-th diagonal entry is $N_{U}(i,i)=\frac{1}{\|U(i,:)\|_{F}}$ for $i\in[n]$, i.e., we have $U_{*}=N_{U}U$. Lemma 3 of Criterion is a distribution-free theoretical result and it shows that there exists a matrix $Y\in\mathbb{R}^{n\times K}_{+}$ such that $U_{*}=YU_{*}(\mathcal{I},:)$, where $Y=N_{M}\Pi\Theta^{-1}(\mathcal{I},\mathcal{I})N^{-1}_{U}(\mathcal{I},\mathcal{I})$ and $N_{M}$ is a diagonal matrix with all diagonal elements being positive. The structure $U_{*}=YU_{*}(\mathcal{I},:)$ is known as Ideal Cone MaoSVM . Set $Z_{*}=UU^{-1}_{*}(\mathcal{I},:)\sqrt{\mathrm{diag}(U_{*}(\mathcal{I},:)\Lambda U^{\prime}_{*}(\mathcal{I},:))}$. When Equation (4) holds such that $P$ has unit-diagonal elements, Equation (8) of Criterion says that $\Pi(i,:)=\frac{Z_{*}(i,:)}{\|Z_{*}(i,:)\|_{1}}$ for $i\in[n]$, i.e., we can recover $\Pi$ from $Z_{*}$ as long as we know the index set $\mathcal{I}$ when $\Omega$ is given, where $Z_{*}$ is directly related to the eigendecomposition of $\Omega$. Since $U_{*}=YU_{*}(\mathcal{I},:)$ is an Ideal Cone structure, the SVM-cone algorithm designed in MaoSVM can be applied to exactly recover $U_{*}(\mathcal{I},:)$ when the inputs are $U_{*}$ and $K$.
@@
Given $\Omega$ and $K$, we provide the Ideal ScD algorithm below, where S denotes SVM, c denotes cone, and D denotes DCMMDF. Input $\Omega,K$. Output: $\Pi$.
  • 1.
@@
  • 2.
    Apply SVM-cone algorithm MaoSVM to $U_{*}$ with $K$ communities to get the index set $\mathcal{I}$.
  • 3.
    Set $Z_{*}=UU^{-1}_{*}(\mathcal{I},:)\sqrt{\mathrm{diag}(U_{*}(\mathcal{I},:)\Lambda U^{\prime}_{*}(\mathcal{I},:))}$.
  • 4.
    Set $\Pi(i,:)=\frac{Z_{*}(i,:)}{\|Z_{*}(i,:)\|_{1}}$ for $i\in[n]$.
Since the SVM-cone algorithm exactly returns $U_{*}(\mathcal{I},:)$, the Ideal ScD exactly recovers $\Pi$, and this also guarantees the identifiability of our DCMMDF.
For the real case when we only know $A$ and $K$, we aim at estimating $\Pi$ when $A$ is generated from any distribution $\mathcal{F}$ satisfying Equation (6) under our DCMMDF. Algorithm 1, which we call ScD, is a natural extension of the Ideal ScD.
@@
3:Let $\tilde{A}=\hat{U}\hat{\Lambda}\hat{U}^{\prime}$ be the leading $K$ eigendecomposition of $A$ such that $\hat{\Lambda}$ contains $A$’s leading $K$ eigenvalues. Let $\hat{U}_{*}$ be the row-normalization of $\hat{U}$ such that $\hat{U}_{*}(i,:)=\frac{\hat{U}(i,:)}{\|\hat{U}(i,:)\|_{F}}$ for $i\in[n]$.
4:Apply SVM-cone algorithm MaoSVM to $\hat{U}_{*}$ with $K$ communities to get the estimated index set $\hat{\mathcal{I}}$.
5:Set $\hat{Z}_{*}=\mathrm{max}(0,\hat{U}\hat{U}^{-1}_{*}(\hat{\mathcal{I}},:)\sqrt{\mathrm{diag}(\hat{U}_{*}(\hat{\mathcal{I}},:)\hat{\Lambda}\hat{U}^{\prime}_{*}(\hat{\mathcal{I}},:))})$.
6:Estimate $\Pi(i,:)$ by setting $\hat{\Pi}_{*}(i,:)=\hat{Z}_{*}(i,:)/\|\hat{Z}_{*}(i,:)\|_{1}$ for $i\in[n]$.
@@
[…]
@@
where $\delta_{K}$ is a threshold and we set it as $\frac{1}{K}$ in this paper. Follow a similar definition of $\hat{\Pi}^{0}$, we can define $\Pi^{0}\in\{0,1\}^{n\times K}$ from $\Pi$. The two binary membership matrices will be used in defining our overlapping weighted modularity.
@@
Because $A$’s expectation is $\Omega$ under the DCMMDF, $\hat{U},\hat{\Lambda},\hat{Z}_{*}$, and $\hat{\Pi}$ are good estimations of $U,\Lambda,Z_{*}$, and $\Pi$, respectively. Again, ScD is the SVM-cone-DCMMSB algorithm of MaoSVM and we re-name the SVM-cone-DCMMSB as ScD to emphasize its distribution-free property. Note that in MaoSVM , the ScD algorithm is only used to estimate $\Pi$ when $A$ is generated from a Bernoulli distribution under DCMM while the ScD used in this article can estimate $\Pi$ when $A$ is generated from any distribution $\mathcal{F}$ as long as Equation (6) holds under our DCMMDF. Furthermore, building spectral algorithms via applications of the eigendecomposition of the adjacency matrix or the Laplacian matrix to fit statistical models for un-weighted networks is a popular way in community detection areas, see rohe2011spectral ; lei2015consistency ; joseph2016impact ; SCORE ; MixedSCORE ; MaoSVM ; mao2020estimating .
@@
The computationally intensive steps of ScD are the leading $K$ eigendecomposition (i.e., SVD) and SVM-cone. The computational cost of SVD is $O(n^{3})$. Because $A$ for real-world networks is usually sparse, applying the power method newman2006modularity , the computation complexity for SVD is $O(Kn^{2})$ SCORE ; MixedSCORE . The complexity of the SVM-cone algorithm is $O(nK^{2})$ chang2011libsvm ; MaoSVM . Since $K\ll n$ in this article, as a result, the total complexity of ScD is $O(Kn^{2})$.

### v3

@@ 2 The Weighted Degree-Corrected Mixed Membership model
Under WDCMM, node $i$’s expectation degree is $\sum_{j=1}^{n}\Omega(i,j)=\theta(i)\Pi(i,:)\sum_{j=1}^{n}P\Pi^{\prime}(j,:)$. We see that if there is a node $\bar{i}$ such that $\Pi(\bar{i},:)=\Pi(i,:)$ (i.e., node $\bar{i}$ has the same community membership as node $i$) for $\bar{i}\neq i$, the expectation degree of node $\bar{i}$ may still differ from that of node $i$ if $\theta(\bar{i})\neq\theta(i)$, and this is the reason we call $\theta(i)$ heterogeneity parameter.
@@
By Proposition A.1 of [33], WDCMM is identifiable when Equations (2)-(4) hold, where Proposition A.1 of [33] has no constraint on $A$’s distribution and it allows $P$ to have negative elements as long as $P$ has unit-diagonals. However, if $P$ has non-unit diagonals, WDCMM is unidentifiable unless all entries of $\theta$ are the same.
@@
In this remark, we compare our WDCMM with some previous models.
  • When $\mathcal{F}$ is Bernoulli distribution, WDCMM reduces to DCMM [33], i.e., WDCMM extends DCMM from un-weighted networks to weighted networks. Given that DCMM includes several well-known models, including MMSB, DCSBM, SBM, and the Erdös-Rényi random graph [61] as special cases, our WDCMM similarly regards these models as special cases. Specifically, by setting $\theta(i)=\sqrt{\rho}$ for all $i\in[n]$, we observe that $\Omega=\rho\Pi P\Pi^{\prime}$, which aligns with the expectation of $A$ as presented in Equation (1) of [36]. Thus, in this particular case, WDCMM degenerates into MMSB. Here, $\rho>0$ is known as the sparsity parameter [24, 36], which governs the network’s sparsity. In this paper, we will theoretically investigate the influence of $\rho$ on the performance of a spectral method by incorporating $\rho$ into our final theoretical bound.
  • When all nodes are pure, WDCMM reduces to DCDFM [53], i.e., WDCMM extends DCDFM from non-overlapping weighted networks to overlapping weighted networks. If we further assume that all entries of $\theta$ are the same, WDCMM reduces to DFM [52].
After designing our model WDCMM, a random adjacency matrix $A$ with community membership $\Pi$ can be generated from WDCMM by the following steps.
@@
After generating the random adjacency matrix $A$ by Steps (a)-(c), in this article, we aim to answer the following questions:
@@
    Can we design an algorithm to estimate $\Pi$ with known $A$ and $K$ when $A$ is generated from arbitrary distribution $\mathcal{F}$ satisfying Equation (6) under our WDCMM?
@@
    Can we design a method to determine the number of communities $K$ for weighted networks generated from WDCMM?
In the next few sections, we will answer the above questions by applying an efficient spectral algorithm to fit WDCMM, building a theoretical guarantee of consistency for the algorithm under WDCMM, and proposing overlapping weighted modularity to infer $K$.
@@ 3 Algorithm
In this section, we aim to answer Q (1) by designing an efficient spectral algorithm to fit WDCMM. Because $\Omega$ has a low dimensional structure with rank $K$ when $\mathrm{rank}(\Pi)=K,\mathrm{rank}(P)=K$, and all diagonal entries of $\Theta$ are positive, to answer Q1, one natural way to design an algorithm to fit WDCMM is from the oracle case with known $\Omega$. A good algorithm fitting WDCMM should satisfy the following nice property: it returns $\Pi$ exactly when using $\Omega$ to replace $A$. Otherwise, if the algorithm can not return $\Pi$ when the input matrix is $\Omega$, it is hard to say that it fits the model well.
@@
To design an algorithm from the oracle case, because $\Omega$ has a low dimensional structure with $K$ nonzero eigenvalues, we let $\Omega=U\Lambda U^{\prime}$ be the eigendecomposition of $\Omega$ such that $U\in\mathbb{R}^{n\times K},\Lambda\in\mathbb{R}^{K\times K}$, and $U^{\prime}U=I_{K}$, where $\Lambda$’s $k$-th diagonal entry is the leading $k$ eigenvalue of $\Omega$ for $k\in[K]$. From now on, we aim at recovering the mixed membership matrix $\Pi$ exactly from $\Omega$’s eigendecomposition $U\Lambda U^{\prime}$. To achieve this goal, we let $U_{*}$ be an $n\times K$ matrix such that its $i$-th row is $U_{*}(i,:)=\frac{U(i,:)}{\|U(i,:)\|_{F}}$ for $i\in[n]$. Let $N_{U}$ be an $n\times n$ diagonal matrix such that its $i$-th diagonal entry is $N_{U}(i,i)=\frac{1}{\|U(i,:)\|_{F}}$ for $i\in[n]$, i.e., we have $U_{*}=N_{U}U$. Lemma 3 of [62] is a distribution-free theoretical result and it shows that there exists a matrix $Y\in\mathbb{R}^{n\times K}_{+}$ such that
$U_{*}=YU_{*}(\mathcal{I},:),$ (7)
where $Y=N_{M}\Pi\Theta^{-1}(\mathcal{I},\mathcal{I})N^{-1}_{U}(\mathcal{I},\mathcal{I})$ and $N_{M}$ is a diagonal matrix with all diagonal elements being positive. Given that no row of $Y$ is 0, $U_{*}(\mathcal{I},:)$ corresponds to $U_{*}$’s $K$ rows, and each row of $U_{*}$ has a unit $l_{2}$ norm, as stated in Problem 1 in [35], the structure $U_{*}=YU_{*}(\mathcal{I},:)$ is known as Ideal Cone [35]. Figure 2 illustrates the Ideal Cone structure inherent in $U_{*}$ when there are three communities. We observe that the mixed rows of $U_{*}$ corresponding to mixed nodes are located at one side of the hyperplane formed by the $K$ pure rows of $U_{*}$ corresponding to pure nodes.
Figure 2: Illustration of the Ideal Cone structure embedded within $U_{*}$ for the case when $K=3$. Here, dots represent the rows of $U_{*}$, while the hyperplane is constituted by the three rows of $U_{*}(\mathcal{I},:)$. Notably, all mixed rows of $U_{*}$ are located on one side of this hyperplane, where we call $U_{*}(i,:)$ a mixed row if the corresponding node $i$ is a mixed node and a pure row otherwise. In this graphical depiction, for every mixed node $i$, we assign its mixed membership as $\Pi(i,1)=r_{1}$, $\Pi(i,2)=r_{2}$, and $\Pi(i,3)=1-r_{1}-r_{2}$, where $r_{1}$ and $r_{2}$ are determined by $\frac{\mathrm{rand}(1)}{2}$, with $\mathrm{rand}(1)$ being a random value drawn from the $\mathrm{Uniform}(0,1)$ distribution. For visualization, these points have been projected and rotated from $\mathbb{R}^{3}$ into $\mathbb{R}^{2}$.
By applying Equation (7), we derive
$U_{*}U^{-1}_{*}(\mathcal{I},:)N_{U}(\mathcal{I},\mathcal{I})\Theta(\mathcal{I},\mathcal{I})=N_{M}\Pi.$ (8)
Recalling that $U_{*}=N_{U}U$, and combing it with Equation (8), we obtain
$UU^{-1}_{*}(\mathcal{I},:)N_{U}(\mathcal{I},\mathcal{I})\Theta(\mathcal{I},\mathcal{I})=N^{-1}_{U}N_{M}\Pi.$ (9)
In Equation (9), given that both $N_{U}$ and $N_{M}$ are diagonal matrices, $\Pi$ can be immediately recovered from $U$ and $\mathcal{I}$ once $\Theta(\mathcal{I},\mathcal{I})$ is known. To determine this term, we note that under the WDCMM model, $\Omega=\Theta\Pi P\Pi^{\prime}\Theta=U\Lambda U^{\prime}$. Consequently, $\Omega(\mathcal{I},\mathcal{I})=\Theta(\mathcal{I},\mathcal{I})\Pi(\mathcal{I},:)P\Pi^{\prime}(\mathcal{I},:)\Theta(\mathcal{I},\mathcal{I})=U(\mathcal{I},:)\Lambda U^{\prime}(\mathcal{I},:)$. Recall that $\Pi(\mathcal{I},:)$ is a $K$-by-$K$ identity matrix, it follows that $\Theta(\mathcal{I},\mathcal{I})P\Theta(\mathcal{I},\mathcal{I})=U(\mathcal{I},:)\Lambda U^{\prime}(\mathcal{I},:)$. Recall that the connectivity matrix $P$ is required to have unit diagonals in Equation (4) for the model’s identifiability, combing this unit-diagonal condition with $\Theta(\mathcal{I},\mathcal{I})P\Theta(\mathcal{I},\mathcal{I})=U(\mathcal{I},:)\Lambda U^{\prime}(\mathcal{I},:)$ immediately yields $\Theta(\mathcal{I},\mathcal{I})=\sqrt{\mathrm{diag}(U(\mathcal{I},:)\Lambda U^{\prime}(\mathcal{I},:))}$. This indicates that a portion of the degree heterogeneity matrix $\Theta$ can be expressed using a portion of the eigendecomposition of $\Omega$. Now, substituting $\sqrt{\mathrm{diag}(U(\mathcal{I},:)\Lambda U^{\prime}(\mathcal{I},:))}$ for $\Theta(\mathcal{I},\mathcal{I})$ in Equation (9), we obtain
$N^{-1}_{U}N_{M}\Pi=UU^{-1}_{*}(\mathcal{I},:)N_{U}(\mathcal{I},\mathcal{I})\sqrt{\mathrm{diag}(U(\mathcal{I},:)\Lambda U^{\prime}(\mathcal{I},:))}.$ (10)
Given that $N_{U}(\mathcal{I},\mathcal{I})$ is a diagonal matrix and $U_{*}=N_{U}U$, it follows that
$N_{U}(\mathcal{I},\mathcal{I})\sqrt{\mathrm{diag}(U(\mathcal{I},:)\Lambda U^{\prime}(\mathcal{I},:))}=\sqrt{\mathrm{diag}(U_{*}(\mathcal{I},:)\Lambda U^{\prime}_{*}(\mathcal{I},:))},$
which simplifies Equation (10) to:
$N^{-1}_{U}N_{M}\Pi=UU^{-1}_{*}(\mathcal{I},:)\sqrt{\mathrm{diag}(U_{*}(\mathcal{I},:)\Lambda U^{\prime}_{*}(\mathcal{I},:))}.$ (11)
Define $Z=UU^{-1}_{*}(\mathcal{I},:)\sqrt{\mathrm{diag}(U_{*}(\mathcal{I},:)\Lambda U^{\prime}_{*}(\mathcal{I},:))}$. By Equation (11), and given that $N^{-1}_{U}N_{M}$ is a diagonal matrix, we immediately deduce
$\Pi(i,:)=\frac{Z(i,:)}{\|Z(i,:)\|_{1}}\mathrm{~for~}i\in[n].$ (12)
Thus, we can recover $\Pi$ from $Z$ using Equation (12) as long as we know the index set $\mathcal{I}$ when $\Omega$ is given, where $Z$ is directly related to the eigendecomposition of $\Omega$ based on Equation (11). In the aforementioned analysis, we observe that:
  • Equations (10)-(12) are valid only when the equality $\Theta(\mathcal{I},\mathcal{I})=\sqrt{\mathrm{diag}(U(\mathcal{I},:)\Lambda U^{\prime}(\mathcal{I},:))}$ holds. This equality, in turn, depends on the unit-diagonal condition imposed on $P$. This further underscores the significance of the unit-diagonal condition on $P$, as it guarantees the identifiability of the WDCMM model. Indeed, if the unit-diagonal condition fails to hold, it would be impossible to accurately recover $\Pi$ from Equation (12), rendering the model unidentifiable.
  • If we disregard $U_{*}$, the row-normalized counterpart of $U$, it becomes impossible to exactly recover $\Pi$ from $U$ due to the degree heterogeneity parameter $\theta$. To mitigate the influence of $\theta$, akin to the methodologies employed in [24, 63, 35], we proceed to normalize each row of $U$ to have unit $l_{2}$ norms.
  • Upon analyzing Equations (7)-(12), readers may observe that we impose no constraints on the elements of the population adjacency matrix $\Omega$, as our focus lies solely on analyzing the structure of $\Omega$’s eigendecomposition. Similarly, readers may notice that apart from the unit-diagonal condition on the connectivity matrix $P$ and the requirement that all elements of $\theta$ must be positive, there are no additional restrictions on $P$ and $\theta$. For instance, the off-diagonal elements of $P$ can be either negative or greater than 1, and the elements of $\theta$ can exceed or fall below 1 in this paper. In essence, we do not impose any constraints on the elements of $\Omega$. These observations underscore the fact that there are no restrictions on the distribution of $A$’s elements, which enhances the generality of our WDCMM model and highlights the advantages of our algorithm, presented later, for weighted networks. In contrast, Mao et al.[35] and Jin et al. [33], who consider the mixed membership estimation task for un-weighted networks, must impose the conditions that all elements of $P$ are non-negative and $0\leq\mathrm{max}_{i,j\in[n]}\Omega(i,j)\leq 1$ in their theoretical analysis.
Since $U_{*}=YU_{*}(\mathcal{I},:)$ is an Ideal Cone structure, the SVM-cone algorithm designed in [35] can be applied to exactly recover $U_{*}(\mathcal{I},:)$ when the inputs are $U_{*}$ and $K$. From the Ideal Cone structure depicted in Figure 2, it is evident that the mixed rows in $U_{*}$, which correspond to mixed nodes, are clearly distinguishable from the pure rows in $U_{*}$, which correspond to pure nodes. The SVM-cone algorithm proposed in [35] is capable of efficiently identifying the pure nodes within the Ideal Cone structure of $U_{*}$.
Remark 3.
Let $\mathcal{I}$ and $\bar{\mathcal{I}}$ be two index sets, where $\mathcal{I}=\{s_{1},s_{2},\ldots,s_{K}\}$ and $\bar{\mathcal{I}}=\{\bar{s}_{1},\bar{s}_{2},\ldots,\bar{s}_{K}\}$ with $s_{k}$ and $\bar{s}_{k}$ being pure nodes in the $k$-th community for $k\in[K]$. We now prove the statement $U_{*}(\mathcal{I},:)=U_{*}(\bar{\mathcal{I}},:)$.
First, we define the matrix $M=\Pi\Theta^{-1}(\mathcal{I},\mathcal{I})U(\mathcal{I},:)$. By Lemma 3 of [62], we know that for any $i\in[n]$, the $(i,i)$-th entry of $N_{M}$ is given by $N_{M}(i,i)=\frac{1}{\|M(i,:)\|_{F}}$. Since $U_{*}=YU_{*}(\mathcal{I},:)=N_{M}\Pi\Theta^{-1}(\mathcal{I},\mathcal{I})N^{-1}_{U}(\mathcal{I},\mathcal{I})U_{*}(\mathcal{I},:)$, we see that for any two distinct nodes $i,\bar{i}$ with the same membership (i.e., $\Pi(i,:)=\Pi(\bar{i},:)$), we have $U_{*}(i,:)=U_{*}(\bar{i},:)$, which leads to $U_{*}(\mathcal{I},:)=U_{*}(\bar{\mathcal{I}},:)$.
This result means that even if the SVM-cone algorithm finds the index set $\bar{\mathcal{I}}$ instead of $\mathcal{I}$, we still obtain the $K\times K$ corner matrix $U_{*}(\mathcal{I},:)$ since $U_{*}(\mathcal{I},:)=U_{*}(\bar{\mathcal{I}},:)$. In other words, for each community, the choice of which pure node $s_{k}$ is picked into the index set $\mathcal{I}$ is not important as long as it is a pure node for $k\in[K]$ for the algorithm to correctly recover the underlying communities.
@@
Given $\Omega$ and $K$, we provide the Ideal ScD algorithm below, where S denotes SVM, c denotes cone, and D denotes distribution-free. Input $\Omega,K$. Output: $\Pi$.
@@
  • Apply SVM-cone algorithm [35] to $U_{*}$ with $K$ communities to get the index set $\mathcal{I}$.
  • Set $Z=UU^{-1}_{*}(\mathcal{I},:)\sqrt{\mathrm{diag}(U_{*}(\mathcal{I},:)\Lambda U^{\prime}_{*}(\mathcal{I},:))}$.
  • Set $\Pi(i,:)=\frac{Z(i,:)}{\|Z(i,:)\|_{1}}$ for $i\in[n]$.
Since the SVM-cone algorithm exactly returns $U_{*}(\mathcal{I},:)$, the Ideal ScD exactly recovers $\Pi$, and this also guarantees the identifiability of our WDCMM.
For the real case when we only know $A$ and $K$, we aim at estimating $\Pi$ when $A$ is generated from any distribution $\mathcal{F}$ satisfying Equation (6) under our WDCMM. Algorithm 1, which we call ScD, is a natural extension of the Ideal ScD.
@@
3:Let $\tilde{A}=\hat{U}\hat{\Lambda}\hat{U}^{\prime}$ be the leading $K$ eigendecomposition of $A$ such that $\hat{\Lambda}$ contains AA’s leading $K$ eigenvalues. Let $S_{+}=\{i\in[n]:\|\hat{U}(i,:)\|_{F}>0\}$ and $\hat{U}^{+}=\hat{U}(S_{+},:)$. Let $\hat{U}_{*}$ be the row-normalization of $\hat{U}^{+}$ such that $\hat{U}_{*}(i,:)=\frac{\hat{U}^{+}(i,:)}{\|\hat{U}^{+}(i,:)\|_{F}}$ for $i\in S_{+}$.
4:Apply SVM-cone algorithm [35] to $\hat{U}_{*}$ with $K$ communities to get the estimated index set $\hat{\mathcal{I}}$.
[…]
6:Estimate $\Pi(i,:)$ by setting $\hat{\Pi}(i,:)=\hat{Z}(i,:)/\|\hat{Z}(i,:)\|_{1}$ for $i\in S_{+}$ and $\hat{\Pi}(i,:)=(\frac{1}{K},\frac{1}{K},\ldots,\frac{1}{K})$ for $i\notin S_{+}$.
Remark 4.
In Algorithm 1, we notice that when the network is extremely sparse, leading to a significant number of zeros in the adjacency matrix $A$, some rows of $\hat{U}$ may turn into zero vectors. To tackle this problem, we employ the strategy presented in Algorithm 2 of [24]. Specifically, we estimate mixed memberships exclusively for the non-zero rows and allocate the mixed memberships of the zero rows equally across all communities.
@@
$\hat{\Pi}^{0}(i,k)=\begin{cases}1&\mbox{when~}\hat{\Pi}(i,k)\geq\delta_{K},\\ 0,&\mbox{otherwise},\end{cases}$
@@
where $\delta_{K}$ is a threshold and we set it as $\frac{1}{K}$ in this paper. Follow a similar definition of $\hat{\Pi}^{0}$, we can define $\Pi^{0}\in\{0,1\}^{n\times K}$ from $\Pi$. The two binary membership matrices will be used in defining our overlapping weighted modularity in Equation (13).
@@
Because $A$’s expectation is $\Omega$ under the WDCMM, $\hat{U},\hat{\Lambda},\hat{Z}$, and $\hat{\Pi}$ are good estimations of $U,\Lambda,Z$, and $\Pi$, respectively. Again, ScD is the SVM-cone-DCMMSB algorithm of [35] and we re-name the SVM-cone-DCMMSB as ScD to emphasize its distribution-free property. Note that in [35], the ScD algorithm is only used to estimate $\Pi$ when $A$ is generated from a Bernoulli distribution under DCMM while the ScD used in this article can estimate $\Pi$ when $A$ is generated from any distribution $\mathcal{F}$ as long as Equation (6) holds under our WDCMM. Furthermore, building spectral algorithms via applications of the eigendecomposition of the adjacency matrix or the Laplacian matrix to fit statistical models for un-weighted networks is a popular way in community detection areas, see [23, 24, 26, 25, 33, 35, 36].
@@
The computationally intensive steps of ScD are the leading $K$ eigendecomposition (i.e., SVD) and SVM-cone. The computational cost of SVD is $O(n^{3})$. Because $A$ for real-world networks is usually sparse, applying the power method [64], the computation complexity for SVD is $O(Kn^{2})$ [25, 33]. The complexity of the SVM-cone algorithm is $O(nK^{2})$ [65, 35]. Since $K\ll n$ in this article, as a result, the total complexity of ScD is $O(Kn^{2})$. In this paper, the notation $a_{1}=O(a_{2})$ indicates that $a_{1}$ and $a_{2}$ are of the same order, meaning $a_{1}$ grows (or decreases) at the same rate as $a_{2}$ increases (or decreases). For instance, $O(Kn^{2})$ signifies that the running time of the ScD algorithm is dominated by terms that scale as the product of $n^{2}$ and $K$, while $a=O(1)$ implies that $a$ is considered to be a constant.
Remark 6.
(Comparison to existing approaches) Here, we compare our ScD algorithm with the SVM-cone-DCMMSB algorithm presented in [35], the Mixed-SCORE algorithm developed in [33], and the OCCAM algorithm introduced in [34].
  • SVM-cone-DCMMSB: As mentioned earlier, our ScD algorithm is essentially the same as the SVM-cone-DCMMSB algorithm due to the following reasons. Our WDCMM model generalizes the DCMM model studied in [35] from un-weighted networks to weighted networks. Both algorithms estimate the mixed membership matrix $\Pi$ using the Ideal Cone structure $U_{*}=YU_{*}(\mathcal{I},:)$, where such structure is initially observed in [35]. However, a key difference lies in the constraints on the adjacency matrix $A$. Our ScD algorithm has no constraints on $A$ since it focuses on weighted networks, whereas the SVM-cone-DCMMSB algorithm considers binary adjacency matrices as it is designed for un-weighted networks.
  • Mixed-SCORE: Recall that our WDCMM model extends the DCMM model introduced in [33] to weighted networks, utilizing the row-wise ratios strategy presented in [25, 33] to eliminate the effect of $\theta$, one can also develop an efficient spectral method to estimate mixed memberships under the WDCMM model. This method is indeed the Mixed-SCORE algorithm developed in [33]. According to Theorem 1.1 in the latest arXiv version of [33], Mixed-SCORE can exactly recover $\Pi$ when the input matrix is the population adjacency matrix $\Omega$. Therefore, the Mixed-SCORE algorithm can be applied to fit our WDCMM model. Developing theoretical guarantees for Mixed-SCORE under WDCMM remains an interesting direction for future work.
  • OCCAM: The OCCAM model introduced in [34] can be viewed as equivalent to the DCMM model [33]. One might naturally wonder whether the OCCAM algorithm designed in [34] could also be applied to fit our WDCMM model. However, this is not the case. The OCCAM algorithm employs the K-medians clustering algorithm to identify cluster centers, which prevents it from exactly recovering the mixed membership matrix $\Pi$ when using the population adjacency matrix $\Omega$ to replace the adjacency matrix $A$ in OCCAM. Consequently, this algorithm is not suitable for fitting our WDCMM model.